# Integral Representations of Three Novel Multiple Zeta Functions for Barnes Type: A Probabilistic Approach

### 02 Jan 2024 v2

@@
(2003, Canad. J. Math., 55, 292β330). The analytically
@@
investigated. Surprisingly, two of them, unlike the previous
results, can extend analytically
to entire functions in the whole complex plane.
Β
@@
Running title: Integral representations of three novel Barnes
@@
Science, Academia Sinica, No.β128, Section 2, Academia Road,
@@
University of Education, No.β1, Jinde Road, Changhua 50007, Taiwan
@@
which plays an important role in analytic number theory (see, e.g.,
Apostal 1976). It is known that each of both zeta functions in (1)
and (2) can extend to a meromorphic function in the whole complex
plane and has a simple pole at where it has residue 1.
@@
the literature (see, e.g., the survey paper by Matsumoto 2010, the
thesis by Onozuka 2014, and the references therein), where
@@
(i) EulerβZagier multiple (Riemann) zeta function (Euler 1776,
Zagier 1994, Hoffman 1992, Zhao 2000, Matsumoto 2002a, Mehta et al.
2016)
(ii) Multiple Hurwitz zeta function (Kamano 2006, Murty and Sinha
2006, Kelliher and Masri 2008, Mehta and Viswanadham 2017)
@@
(iii) Barnes multiple zeta function (of one variable) (Barnes 1901, 1904, Choi and Quine 1992, Ruijsenaars 2000, Matsumoto 2002a, 2010)
@@
, see Shintani (1977, 1978) and Mellin (1902, 1904) for
the case of one variable, and Matsumoto (2002a, 2002b, 2003a, 2003b,
2005) for the case of several variables. The analytically continued
properties of the above-mentioned multiple zeta functions have been
investigated. For example, the Barnes multiple zeta function
extends to a meromorphic function in and has
simple poles only at (see Barnes 1904, Ruijsenaars
2000 or Matsumoto 2010). The main tools of analytic continuation of
the zeta functions include integral representation, MellinβBarnes
integral formula, EulerβMaclaurin summation, translation formula
(identity), generalized function, etc. For the recent development on
the multiple zeta functions, see the monograph by Zhao (2016).
In this paper, we will introduce the following three variants (a) β
@@
Shintani double zeta functions of the form (Shintani 1975, Kim et
al. 2022)
@@
Then, we define
@@
distributions given by Pitman and Yor (2003), and investigate their
analytically continued properties including the set of
@@ 2. Lemmas
the basic properties of the gamma function (see, e.g., Apostal 1976,
Chapter 12). Lemma 2 is about Liouvilleβs extension of Dirichlet
integral theorem (see, e.g., Sharma 2005, Chapter 2, or Gibson
1931, p.β492). The crucial Lemma 3 modifies slightly
Cheng and Huβs (2022) Lemma 2.1. The
@@
investigated by Pitman and Yor (2003). On the other hand, for the
derivations of the hyperbolic density functions in Lemma 5, see
Cheng and Hu (2022) as well as Pitman and Yor (2003).
@@
Lemma 2 (LiouvilleβDirichlet integral). Let
@@
Then we have, by Cauchyβs theorem,
@@
Moreraβs theorem completes the proof.
@@
notation ββ means equality in distribution.
@@
Corollary 2 (Cheng and Hu 2022). The Hurwitz
zeta function has the integral representation of the
form
@@
Corollary 5 (Cheng and Hu 2022). The zeta
function in has the integral representation
@@
Corollary 6 (Cheng and Hu 2022). The zeta
function in has the integral representation
of the form
@@
Corollary 9 (Cheng and Hu 2022). The zeta
function in has the integral
representation
@@ 4. Proofs of the main results
series on the RHS of (3) converges absolutely provided To see this, note that by Stirlingβs formula, or more
@@
Now, by (29), (30), the above claim and Fubiniβs theorem, we can
@@
Inserting (32) in (31) yields, by (29) and Fubiniβs theorem again,
@@
Inserting (36) in (35) yields, by (29) and Fubiniβs theorem again,
@@
Then, by (29) and Fubiniβs theorem, we can rewrite the expectation
@@
Then it follows from (38)β(41) and the definition of the random

### 31 Mar 2026 v3

@@
(Canad. J. Math., 55 (2003), 292–330). The analytically
@@
investigated. The first one generalizes the well-known Barnes
multiple zeta function, while the second and the third, unlike the
previous results, can extend analytically to entire functions in the
whole complex plane. These are somewhat surprising results.
@@
Running title: Three novel Barnes
@@
Science, Academia Sinica, No. 128, Section 2, Academia Road,
@@
University of Education, No. 1, Jinde Road, Changhua 50007, Taiwan
@@
which plays an important role in analytic number theory (see, e.g., Apostal
[1]).
It is known that each of both zeta functions in (1) and (2) can
extend to a meromorphic function in the whole complex plane and
has a simple pole at where it has residue 1.
@@
the literature (see, e.g., the survey paper by Matsumoto
[17], the thesis by Onozuka [23], and the
references therein), where
@@
(i) Euler–Zagier multiple (Riemann) zeta function ([6, 8, 12, 18, 30, 31])
(ii) Multiple Hurwitz zeta function ([9, 10, 19, 22])
@@
(iii) Barnes multiple zeta function (of one variable) ([2, 3, 5, 12, 17, 25])
@@
, see [20, 21, 28, 29] for the
case of one variable, and [12]–[16] for the
case of several variables. The analytically continued properties of
the above-mentioned multiple zeta functions have been investigated.
For example, the Barnes multiple zeta function extends
to a meromorphic function in and has simple poles only at
(see [3, 17, 25]). The main
tools of analytic continuation of the zeta functions include
integral representation, Mellin–Barnes integral formula,
Euler–Maclaurin summation, translation formula (identity),
generalized function, etc. For the recent development on the
multiple zeta functions, see the monograph [32].
In this paper, we will introduce the following three variants (a) –
@@
Shintani double zeta functions of the form ([11, 27])
@@
Then, we define (see Remark 1 below for the names of moment functions)
@@
distributions given by Pitman and Yor [24], and investigate
their analytically continued properties including the set of
@@
the basic properties of the gamma function (see, e.g., Apostal [1, Chapter
12]). Lemma 2 is about Liouville’s extension of Dirichlet
integral theorem (see, e.g., Sharma [26, Chapter 2], or Gibson
[7, page 492]). The crucial Lemma 3 modifies slightly
Cheng and Hu’s [4] Lemma 2.1. The
@@
investigated by Pitman and Yor [24]. On the other hand, for
the derivations of the hyperbolic density functions in Lemma 5, see
[4, 24].
@@
Lemma 2 (Liouville–Dirichlet integral). Let
@@
Then we have, by applying first Fubini’s theorem and then
Cauchy’s theorem,
@@
Morera’s theorem completes the proof.
@@
notation “” means equality in distribution.
@@
Corollary 2 ([4]). The Hurwitz zeta
function has the integral representation
@@
Corollary 5 ([4]). The zeta function
in has the integral representation
@@
Corollary 6 ([4]). The zeta function
in has the integral representation
@@
Corollary 9 ([4]). The zeta function
in has the integral representation
@@ 4. Proofs of the main results
series on the RHS of (3) converges absolutely provided To see this, note that by Stirling’s formula, or more
@@
Now, by (29), (30), the above claim and Fubini’s theorem, we can
@@
Inserting (32) in (31) yields, by (29) and Fubini’s theorem again,
@@
Inserting (36) in (35) yields, by (29) and Fubini’s theorem again,
@@
Then, by (29) and Fubini’s theorem, we can rewrite the expectation
@@
Then it follows from (38)–(40) and the definition of the random